# HydroGEM: A Self Supervised Zero Shot Hybrid TCN Transformer Foundation Model for Continental Scale Streamflow Quality Control

### v1

@@ 3.1.1 USGS Streamflow Archive
We assembled a continental-scale dataset from 5,321 USGS streamgaging stations via the National Water Information System (NWIS) [usgs2024nwis], covering January 2000 through December 2024. For each station, we selected the best 10-year continuous period with high-quality paired discharge ($Q$) and stage ($H$) observations. All records carry USGS qualification code ‘A’ (approved), distinguishing them from provisional data subject to revision. Instantaneous readings (typically 15-minute intervals) were aggregated to hourly resolution via arithmetic means for consistent temporal scale.
@@ 3.1.2 Spatial and Hydrologic Coverage
This spatial coverage ensures exposure to principal North American flow generation mechanisms: snowmelt-dominated montane regimes, rainfall-dominated humid systems, mixed snow-rain transitional climates, ephemeral arid flows, and regulated downstream conditions [chow1988applied]. Such hydrologic diversity is critical for developing transferable representations that generalize beyond specific flow regimes.
@@ 3.1.3 Quality Control and Preprocessing
Gap filling uses hierarchical strategies: linear interpolation for gaps $\leq 6$ hours, exponential recession models [vogel1992baseflow, tallaksen1995hydrograph] for 6-24 hour gaps, and exclusion for gaps >24>24 hours. This protocol retains 94.7% of potential sequences, with 89.3% containing no interpolation. Windows intersecting flagged spans are dropped, eliminating information leakage.
@@ 3.1.4 Dataset Partitioning
To assess generalization, we assign entire sites to a single partition, avoiding leakage of site-specific signatures [kratzert2019towards]. The split allocates 70% training (3724 sites), 15% validation (798 sites), and 15% test (799 sites), preserving geographic coverage and proportional hydrologic region representation.
Continuous hourly series are segmented into 576-hour windows (24 days), capturing storm hydrographs, weekly patterns, and diurnal fluctuations while remaining tractable for transformers [vaswani2017attention]. Asymmetric stride—48 hours for training (dense overlap maximizing data utilization), 192 hours for validation/test (reducing autocorrelation)—produces 6.03M training sequences, 0.32M validation, and 0.33M test sequences (Table 2).
@@ 3.2.1 Feature Engineering
Static descriptors $\{\phi,\lambda,A_{d},z_{g}\}$ provide coarse hydrograph controls through latitude (climatic gradients), longitude (monsoon influence), drainage area (flashiness), and elevation (temperature lapse rates). Dynamic variables $\{Q,H\}$ constitute the core physical state, with stage encoding local geometry and backwater effects [nittrouer2013backwater] invisible in discharge alone. Derived features include scale embeddings returning absolute variability information lost during standardization, cross-site ranks stabilizing learning across magnitude orders, and monthly anomalies expressing seasonal departures.
@@ 3.2.2 Hierarchical Normalization
Tier 1 - Logarithmic stabilization: Apply $Q^{(1)}=\ln(Q+\epsilon)$, $H^{(1)}=\ln(H+\epsilon)$ for approximately log-normal variables [sangal1970lognormal], linearizing rating curves and stabilizing variance.
@@ 3.4.1 Synthetic Test Set Design
Geographic separation: Zero site overlap eliminates memorization of station-specific patterns. Mathematical separation: Each test anomaly type implements 3-4 equation variants (e.g., drift: linear, exponential, sigmoid, polynomial) producing similar visual patterns through different mechanisms. Temporal separation: Structured duration regimes (micro/meso/macro) spanning 3-520 hours, with macro extended to 520h based on expert feedback regarding ice persistence [santeford1986ice]. Parameter separation: Coverage and severity anti-overlap the training sweet spot, testing distribution boundaries.
@@ 4.1 Two-Stage Training Framework
HydroGEM uses a two-stage training approach that decouples representation learning from task-specific optimization, following recent advances in foundation model development [devlin2019bert, He2022]. This architectural design addresses the challenge of learning robust hydrological representations from limited labeled anomaly data while maintaining generalization across diverse monitoring sites and flow regimes.
@@
This decoupled approach provides several advantages over end-to-end anomaly detection architectures. The self-supervised pretraining exploits massive volumes of unlabeled data that would be prohibitively expensive to manually annotate, learning subtle hydrological patterns that might not be captured from limited labeled examples. The pretrained representations show superior transfer learning capabilities [TransferLearning2021], allowing detection of new anomaly types not explicitly included in the fine-tuning dataset. Furthermore, the modular design permits independent optimization of reconstruction quality and detection performance, facilitating ablation studies and architectural improvements without complete retraining.
@@ 4.2.1 Backbone Network Architecture
The TCN output is projected from 128 to 256 dimensions and processed through 4 transformer layers with 8-head attention. We use Cosine Retention Attention with learnable temporal decay [Mongaras2024] to prevent attention collapse in long sequences. Sliding window attention (256 positions) reduces complexity from $O(T^{2})$ to $O(T\cdot W)$. The transformer contributes 9.8 million parameters.
@@ 4.2.2 Masking Strategy and Objectives
The distinction between missing and corrupted data is critical. Missing data (Stage 1) requires the model to leverage contextual information and physical relationships to infer likely values, teaching it the underlying structure of hydrological systems. Corrupted data (Stage 2) contains incorrect values that must be identified and corrected, a task made possible by the model’s pretrained understanding of what constitutes normal behavior. This two-stage approach has proven successful in vision [He2022] and language understanding [devlin2019bert], and we adapt it here for hydrological quality control.
We implement four masking patterns calibrated to the typical data gaps encountered in operational monitoring, informed by masking strategies in self-supervised learning [Wettig2023]:
@@
  • Feature masking (10% probability): Masks either discharge (70%) or stage (30%) for 24-168 hours, forcing the model to learn rating curve relationships [Kennedy1984] and cross-variable dependencies.
@@ 4.2.3 Pretraining Loss Functions
The backbone trains for 20 epochs using AdamW optimization [Loshchilov2017] with OneCycleLR scheduling (peak learning rate $5\times 10^{-4}$), gradient clipping (norm 1.0), and early stopping (patience 7 epochs). Training requires approximately 48 GPU-hours on NVIDIA A100 hardware.
@@ 4.3.1 Detection Head Architecture
The detection head operates entirely on observable quantities without requiring ground-truth clean data during inference [Angiulli2023, Gong2019]. It computes an 11-dimensional feature vector from the relationship between potentially corrupted observations and backbone reconstructions:
@@ 5.2 Evaluation Metrics
For the Canadian case study, reconstructions are evaluated more qualitatively. Weak labels are derived from operational corrections, and correction strategies, especially for ice affected periods where discharge is often pushed very low while stage remains relatively high, differ from the USGS corrections used to train HydroGEM. Official guidance notes that stage affected by backwater from ice may accurately represent local conditions but may not be suitable for discharge computation, and operational products explicitly warn that ice can raise stage for a given flow [usgs_wdr_docs, eccc_data_manual, eccc_ice_warning]. In our sample many Environment and Climate Change Canada winter corrections reduce discharge much more aggressively than the USGS style corrections seen in training, particularly during ice control periods where stage remains elevated. As a result, in the Canadian evaluation we treat reconstruction error as a secondary diagnostic rather than a primary performance metric and place emphasis on detection metrics.
@@ 5.4 Evaluation Protocols
For cross national evaluation we apply HydroGEM without retraining to 100 Canadian stations from Environment and Climate Change Canada. Weak labels are derived from human corrections in the operational record. These labels reflect editorial decisions and local correction practice, so we report F1 and related detection metrics over a range of labeling thresholds defined by the relative magnitude of human edits, with full curves in the appendix. This provides a sensitivity view of how performance changes as the notion of meaningful correction becomes stricter. Given the documented differences between USGS and Environment and Climate Change Canada correction strategies in ice affected periods [usgs_wdr_docs, eccc_data_manual, eccc_ice_warning], we interpret reconstruction errors here as supportive evidence rather than a primary outcome and focus our conclusions on detection performance.
@@ 6.2.2 Representative Detection Example
Figure 7: Detection example for Canadian station 05AD028 (F1 = 0.859). Blue and red shading indicate periods where ground truth and HydroGEM identify anomalies, respectively. The model preserves clean observations (late March) while flagging data quality issues. Reconstruction divergence during ice-affected periods (early March) reflects differences between USGS training corrections and ECCC operational practices [Turcotte2013, Beltaos2013].
@@ 6.2.3 Detection Behavior Characterization
Beyond aggregate metrics, we analyzed whether HydroGEM learned physically meaningful patterns that transfer to Canadian conditions. Figure 8 summarizes detection behavior across seasons, correction magnitudes, and anomaly durations.
@@
We compared HydroGEM’s flag rate against the human correction rate across seasons (Figure 8A). Both exhibit consistent seasonal structure: winter shows the highest activity (HydroGEM: 53.6%, human: 68.2%), followed by spring (50.6%, 59.5%), fall (42.9%, 56.4%), and summer (39.3%, 42.4%). The winter peak reflects well-documented challenges in cold-region hydrometry: ice-affected flow produces backwater conditions violating standard rating curves, anchor ice and frazil ice cause sensor interference, and instrument freezing creates data gaps [WMO2010a, Turcotte2013, Beltaos2013]. That HydroGEM—trained on predominantly ice-free USGS stations—identifies these signatures in Canadian data suggests the model learned generalizable representations of sensor malfunction and rating curve violation.
@@
A concern with learned detectors is whether they simply flag large deviations while missing subtle corrections. Table 8 stratifies detection recall by correction magnitude. HydroGEM maintains consistent recall (0.48–0.56) across corrections ranging from minor adjustments (1–5%) to major revisions (50–100%), indicating the model learned anomaly patterns—temporal dynamics, channel inconsistencies, physical constraint violations—rather than relying on deviation magnitude as a proxy.
Table 8: Detection recall by correction magnitude. Consistent performance across magnitude ranges indicates pattern-based rather than magnitude-based detection.
Correction Magnitude Timesteps Detected Recall
1–5% 209,223 114,903 0.52
5–10% 54,174 32,297 0.49
10–25% 69,868 36,054 0.48
25–50% 63,022 33,030 0.52
50–100% 141,958 77,670 0.56
Temporal scale coverage.
Operational data quality issues span diverse temporal scales. Analysis of contiguous anomaly segments reveals this heterogeneity: 48% of correction events span fewer than 6 hours (sensor spikes, transient failures), 29% span 6–48 hours (storm-event artifacts, equipment malfunctions), and 23% exceed 48 hours (extended ice cover, prolonged sensor failures). HydroGEM’s training on synthetic anomalies spanning multiple temporal scales allows detection across this full spectrum (Figure 8C).
Figure 8: Detection behavior on Canadian zero-shot evaluation. (A) Seasonal comparison of HydroGEM flag rate versus human correction rate, showing aligned structure with winter peaks. (B) Per-site agreement between model and human corrections. (C) Distribution of anomaly event durations. (D) Detection recall by correction magnitude.
@@ 7.1 Interpretation of Key Findings
Zero-shot transfer to Canadian stations shows that learned representations capture universal principles of streamflow monitoring rather than USGS-specific artifacts. The alignment between HydroGEM’s detection patterns and operational correction rates across seasons is particularly striking. The winter peak in both model flagging and human corrections, corresponding to documented ice effects [Turcotte2013, Beltaos2013], emerges despite training predominantly on ice-free stations. This suggests the model learned to identify fundamental signatures of sensor malfunction and rating curve violation that manifest similarly across different hydrological contexts. The ability to generalize across national boundaries, instrumentation protocols, and climatic regimes validates the foundation model approach for operational hydrology.
@@ 7.2 Limitations and Operational Considerations
Reconstruction divergence in ice-affected periods represents the most significant operational challenge. While HydroGEM successfully detects ice-related anomalies with high recall, reconstructions during these periods often diverge from operational corrections, particularly for Canadian stations. This divergence reflects fundamental differences in correction philosophy between agencies. USGS guidance emphasizes maintaining physical plausibility of both discharge and stage, typically applying moderate adjustments that preserve the coupled relationship. In contrast, ECCC practice often aggressively reduces discharge to near-zero while preserving elevated stage readings during ice backwater, acknowledging that stage may accurately represent local conditions even when unsuitable for discharge computation [eccc_ice_warning]. This is not a detection failure but rather highlights that “correct” values depend on operational context, downstream use cases, and agency-specific protocols. For deployment, we strongly recommend treating all reconstructions as suggestions requiring expert review, particularly during known ice periods.
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. EAR 2012123. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Any use of trade, firm, or product names is for descriptive purposes only and does not imply endorsement by the U.S. Government.
The work was also supported by the University of Vermont College of Engineering and Mathematical Sciences through the REU program.
We acknowledge the US Geological Survey (USGS) and Environment and Climate Change Canada (ECCC) for making the hydrometric datasets publicly available. We also thank David Baude, Research Analyst at the Water Resources Institute, University of Vermont, for his expert review and valuable feedback on the synthetic anomaly taxonomy.
@@ A.0.1 Tier 1: Logarithmic Stabilization
For approximately log-normal hydrological variables [sangal1970lognormal], we apply:
@@
Natural channel rating curves follow power-law form $Q=a(H-H_{0})^{b}$ [chow1988applied]. In log-space:

### v2

@@ 2.1.1 USGS Streamflow Archive
We assembled a continental-scale dataset from 5,321 USGS streamgaging stations via the National Water Information System (NWIS) [55], covering January 2000 through December 2024. For each station, we selected the best 10-year continuous period with high-quality paired discharge ($Q$) and stage ($H$) observations. All records carry USGS qualification code ‘A’ (approved), distinguishing them from provisional data subject to revision. Instantaneous readings (typically 15-minute intervals) were aggregated to hourly resolution via arithmetic means for consistent temporal scale.
@@ 2.1.2 Spatial and Hydrologic Coverage
This spatial coverage ensures exposure to principal North American flow generation mechanisms: snowmelt-dominated montane regimes, rainfall-dominated humid systems, mixed snow-rain transitional climates, ephemeral arid flows, and regulated downstream conditions [9]. Such hydrologic diversity is critical for developing transferable representations that generalize beyond specific flow regimes.
@@ 2.1.3 Quality Control and Preprocessing
Gap filling uses hierarchical strategies: linear interpolation for gaps $\leq 6$ hours, exponential recession models [57, 52] for 6-24 hour gaps, and exclusion for gaps >24>24 hours. This protocol retains 94.7% of potential sequences, with 89.3% containing no interpolation. Windows intersecting flagged spans are dropped, eliminating information leakage.
@@ 2.1.4 Dataset Partitioning
To assess generalization, we assign entire sites to a single partition, avoiding leakage of site-specific signatures [27]. The split allocates 70% training (3724 sites), 15% validation (798 sites), and 15% test (799 sites), preserving geographic coverage and proportional hydrologic region representation.
Continuous hourly series are segmented into 576-hour windows (24 days), capturing storm hydrographs, weekly patterns, and diurnal fluctuations while remaining tractable for transformers [56]. Asymmetric stride—48 hours for training (dense overlap maximizing data utilization), 192 hours for validation/test (reducing autocorrelation)—produces 6.03M training sequences, 0.32M validation, and 0.33M test sequences (Table 2).
@@ 2.2.1 Feature Engineering
Static descriptors $\{\phi,\lambda,A_{d},z_{g}\}$ provide coarse hydrograph controls through latitude (e.g., climatic gradients), longitude (e.g., continental position), drainage area (e.g., flashiness), and elevation (e.g., temperature lapse rates). Dynamic variables $\{Q,H\}$ constitute the core physical state, with stage encoding local geometry and backwater effects [39] invisible in discharge alone. Derived features include scale embeddings returning absolute variability information lost during standardization, cross-site ranks stabilizing learning across magnitude orders, and monthly anomalies expressing seasonal departures.
@@ 2.2.2 Hierarchical Normalization
Tier 1 - Logarithmic stabilization: Apply $Q^{(1)}=\ln(Q+\epsilon)$, $H^{(1)}=\ln(H+\epsilon)$ for approximately log-normal variables [45], linearizing rating curves and stabilizing variance.
@@ 2.4.1 Synthetic Test Set Design
Geographic separation: Zero site overlap eliminates memorization of station-specific patterns. Mathematical separation: Each test anomaly type implements 3-4 equation variants (e.g., drift: linear, exponential, sigmoid, polynomial) producing similar visual patterns through different mechanisms. Temporal separation: Structured duration regimes (micro/meso/macro) spanning 3-520 hours, with macro extended to 520h based on expert feedback regarding ice persistence [46]. Parameter separation: Coverage and severity anti-overlap the training sweet spot, testing distribution boundaries.
@@ 3.1 Two-Stage Training Framework
HydroGEM uses a two-stage training approach that decouples representation learning from task-specific optimization, following recent advances in foundation model development [10, 20]. This architectural design addresses the challenge of learning robust hydrological representations from limited labeled anomaly data while maintaining generalization across diverse monitoring sites and flow regimes.
@@
This decoupled approach provides several advantages over end-to-end anomaly detection architectures. The self-supervised pretraining exploits massive volumes of unlabeled data that would be prohibitively expensive to manually annotate, learning subtle hydrological patterns that might not be captured from limited labeled examples. The pretrained representations show superior transfer learning capabilities [60], allowing detection of new anomaly types not explicitly included in the fine-tuning dataset. Furthermore, the modular design permits independent optimization of reconstruction quality and detection performance, facilitating ablation studies and architectural improvements without complete retraining.
@@ 3.2.1 Backbone Network Architecture
The TCN output is projected from 128 to 256 dimensions and processed through 4 transformer layers with 8-head attention. We use Cosine Retention Attention with learnable temporal decay [36] to prevent attention collapse in long sequences. Sliding window attention (256 positions) reduces complexity from $O(T^{2})$ to $O(T\cdot W)$. The transformer contributes 9.8 million parameters.
@@ 3.2.2 Masking Strategy and Objectives
The distinction between missing and corrupted data is critical. Missing data (Stage 1) requires the model to leverage contextual information and physical relationships to infer likely values, teaching it the underlying structure of hydrological systems. Corrupted data (Stage 2) contains incorrect values that must be identified and corrected, a task made possible by the model’s pretrained understanding of what constitutes normal behavior. This two-stage approach has proven successful in vision [20] and language understanding [10], and we adapt it here for hydrological quality control.
We implement four masking patterns calibrated to the typical data gaps encountered in operational monitoring, informed by masking strategies in self-supervised learning [59]:
@@
  • Feature masking (10% probability): Masks either discharge (70%) or stage (30%) for 24-168 hours, forcing the model to learn rating curve relationships [25] and cross-variable dependencies.
@@ 3.2.3 Pretraining Loss Functions
The backbone trains for 20 epochs using AdamW optimization [33] with OneCycleLR scheduling (peak learning rate $5\times 10^{-4}$), gradient clipping (norm 1.0), and early stopping (patience 7 epochs). Training requires approximately 48 GPU-hours on NVIDIA A100 hardware.
@@ 3.3.1 Detection Head Architecture
The detection head operates entirely on observable quantities without requiring ground-truth clean data during inference [3, 16]. It computes an 11-dimensional feature vector from the relationship between potentially corrupted observations and backbone reconstructions:
@@ 4.2 Evaluation Metrics
For the Canadian case study, reconstructions are evaluated more qualitatively. Weak labels are derived from operational corrections, and correction strategies, especially for ice affected periods where discharge is often pushed very low while stage remains relatively high, differ from the USGS corrections used to train HydroGEM. Official guidance notes that stage affected by backwater from ice may accurately represent local conditions but may not be suitable for discharge computation, and operational products explicitly warn that ice can raise stage for a given flow [41, 44, 18]. In our sample many Environment and Climate Change Canada winter corrections reduce discharge much more aggressively than the USGS style corrections seen in training, particularly during ice control periods where stage remains elevated. As a result, in the Canadian evaluation we treat reconstruction error as a secondary diagnostic rather than a primary performance metric and place emphasis on detection metrics.
@@ 4.4 Evaluation Protocols
For cross national evaluation we apply HydroGEM without retraining to 100 Canadian stations from Environment and Climate Change Canada. Weak labels are derived from human corrections in the operational record. These labels reflect editorial decisions and local correction practice, so we report F1 and related detection metrics over a range of labeling thresholds defined by the relative magnitude of human edits, with full curves in the appendix. This provides a sensitivity view of how performance changes as the notion of meaningful correction becomes stricter. Given the documented differences between USGS and Environment and Climate Change Canada correction strategies in ice affected periods [41, 44, 18], we interpret reconstruction errors here as supportive evidence rather than a primary outcome and focus our conclusions on detection performance.
@@ 5.2.2 Representative Detection Example
Figure 7: Detection example for Canadian station 05AD028 (F1 = 0.859). Blue and red shading indicate periods where ground truth and HydroGEM identify anomalies, respectively. The model preserves clean observations (late March) while flagging data quality issues. Reconstruction divergence during ice-affected periods (early March) reflects differences between USGS training corrections and ECCC operational practices [54, 6].
@@ 5.2.3 Detection Behavior Characterization
Beyond aggregate metrics, we analyzed whether HydroGEM learned physically meaningful patterns that transfer to Canadian conditions. Figure 8 summarizes detection behavior across seasons and correction magnitudes.
@@
We compared HydroGEM’s flag rate against the human correction rate across seasons (Figure 8A). Both exhibit consistent seasonal structure: winter shows the highest activity (HydroGEM: 54%, human: 68%), followed by spring (51%, 60%), fall (43%, 56%), and summer (39%, 42%). The winter peak reflects well-documented challenges in cold-region hydrometry: ice-affected flow produces backwater conditions violating standard rating curves, anchor ice and frazil ice cause sensor interference, and instrument freezing creates data gaps [34, 54, 6]. That HydroGEM—trained on predominantly ice-free USGS stations—identifies these signatures in Canadian data suggests the model learned generalizable representations of sensor malfunction and rating curve violation.
@@
A concern with learned detectors is whether they simply flag large deviations while missing subtle corrections. Figure 8B stratifies detection recall by correction magnitude. HydroGEM maintains consistent recall (0.48–0.56) across corrections ranging from minor adjustments (1–5%) to major revisions (50–100%), indicating the model learned anomaly patterns—temporal dynamics, channel inconsistencies, physical constraint violations—rather than relying on deviation magnitude as a proxy.
Figure 8: Detection behavior on Canadian zero-shot evaluation. (A) Seasonal comparison of HydroGEM flag rate versus human correction rate, showing aligned patterns with winter peaks. (B) Detection recall by correction magnitude, demonstrating consistent performance across small to large corrections.
@@ 6.1 Interpretation of Key Findings
Zero-shot transfer to Canadian stations shows that learned representations capture universal principles of streamflow monitoring rather than USGS-specific artifacts. The alignment between HydroGEM’s detection patterns and operational correction rates across seasons is particularly striking. The winter peak in both model flagging and human corrections, corresponding to documented ice effects [54, 6], emerges despite training predominantly on ice-free stations. This suggests the model learned to identify fundamental signatures of sensor malfunction and rating curve violation that manifest similarly across different hydrological contexts. The ability to generalize across national boundaries, instrumentation protocols, and climatic regimes validates the foundation model approach for operational hydrology.
@@ 6.2 Limitations and Operational Considerations
Reconstruction divergence in ice-affected periods represents the most significant operational challenge. While HydroGEM successfully detects ice-related anomalies with high recall, reconstructions during these periods often diverge from operational corrections, particularly for Canadian stations. This divergence reflects fundamental differences in correction philosophy between agencies. USGS guidance emphasizes maintaining physical plausibility of both discharge and stage, typically applying moderate adjustments that preserve the coupled relationship. In contrast, ECCC practice often aggressively reduces discharge to near-zero while preserving elevated stage readings during ice backwater, acknowledging that stage may accurately represent local conditions even when unsuitable for discharge computation [18]. This is not a detection failure but rather highlights that “correct” values depend on operational context, downstream use cases, and agency-specific protocols. For deployment, we strongly recommend treating all reconstructions as suggestions requiring expert review, particularly during known ice periods.
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No. EAR 2012123. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Any use of trade, firm, or product names is for descriptive purposes only and does not imply endorsement by the U.S. Government. The work was also supported by the University of Vermont College of Engineering and Mathematical Sciences through the REU program. The authors would like to thank the US Geological Survey (USGS) for offering the domain expertise that was crucial to identify the peak anomaly types that are of practical importance. *** Ijaz, I think we should acknowledge USGS and ECCC for the datasets made available for us. *** The authors acknowledge the National Artificial Intelligence Research Resource (NAIRR) Pilot and the Texas Advanced Computing Center for contributing to the research by providing GPU computing resources under Award No. NAIRR240491.
@@ A.0.1 Tier 1: Logarithmic Stabilization
For approximately log-normal hydrological variables [45], we apply:
@@
Natural channel rating curves follow power-law form $Q=a(H-H_{0})^{b}$ [9]. In log-space: